OpenStudy (anonymous):

given f(x)=2/(x^2-x), simplify into a single fraction that follows the expression: f(x+h)-f(x).

OpenStudy (anonymous):

\[f(x)=\frac{2}{x^{2}-x}=\frac{2}{x(x-1)}=\frac{2x-2x+2}{x(x-1}=\frac{2x}{x(x-1)}-\frac{2x-2}{x(x-1)}\]\[=\frac{2}{x-1}-\frac{2}{x}\]